# SAFES: Sequential Privacy and Fairness Enhancing Data Synthesis for Responsible AI

### v1

@@ 4.3 Fairness-aware data transformation
for all $y$ and groups $g_{1},g_{2}\in G$. $\eta$ controls the trade-off between enforcing group fairness and maintaining statistical relationships in the input dataset; different $\eta$ can be used for different groups or response values. We denote $\eta_{y,g_{1},g_{2}}$ for different $\{y,g_{1},g_{2}\}$ collectively by $\boldsymbol{\eta}$. Enforcing group fairness via equation (8) risks changing certain individuals unrealistically. To balance group and individual fairness, TOT limits the conditional expected distortion. Let $\phi:(\mathcal{X}\times\mathcal{Y})^{2}\rightarrow\mathbb{R}$ be a nonnegative distortion function, where $\mathcal{X}$ and $\mathcal{Y}$ are the domains of $X$ and $Y$. TOT requires
@@
$P\left(\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=1|G=\mathbf{g},X=\mathbf{x},Y=y\right)\leq c_{\mathbf{g},\mathbf{x},y}.$ (10)
Similar to $\eta$, different values of $c_{\mathbf{g},\mathbf{x},y}$ can be used for different groups or response values. We denote the bounds $c_{\mathbf{g},\mathbf{x},y}$ collectively as $\mathbf{c}$. The distortion function is also customizable. This customization should be based on the context of the problem, and there is not necessarily a single “best” function for a given case. For example, a mapping that modifies an individual’s age by several decades should receive a high distortion value $\phi$, while a mapping that does not change an individual’s age would receive a 00 distortion value. Finally, to ensure the transformation does not drastically modify the distribution of the input dataset, the objective function minimizes the total variation distance (TVD), which, for discrete attributes, is
@@
Equations (8), (9), and (11), plus the constraint that $q_{\tilde{X},\tilde{Y}|D}$ is a valid distribution, give the optimization
@@
$J(q_{\tilde{Y}|G}(y|g_{1}),q_{\tilde{Y}|G}(y|g_{2}))\leq\eta_{y,g_{1},g_{2}},$ (13)
$\mathbb{E}\left[\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})|G=\mathbf{g},X=\mathbf{x},Y=y\right]\leq c_{\mathbf{g},\mathbf{x},y},$
which can be solved via a standard convex solver like the embedded conic solver (ECOS) (Domahidi et al.,, 2013) available in the CVXPY library (Diamond and Boyd,, 2016).
@@ 5 Experiments
We run two experiments using the SAFES procedure on the Adult (also known as the 1994 US “Census Income”) (Becker and Kohavi,, 1996) and the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) (ProPublica,, 2016) datasets. Each of these datasets are publicly available for download. We evaluate the three-way privacy-fairness-utility trade-offs in SAFES against data synthesis or transformation methods that focus on either privacy preservation or fairness enhance, but not both, as well as against a baseline with no privacy or fairness enhancing procedure applied.
@@ 5.1 Data
The Adult dataset is a subset of US 1994 Census income data on 48,842 individuals. The dataset is frequently used in both privacy and fairness literature due to the inclusion of sensitive variables (e.g., income) and the encoding of real-world discrimination (e.g., pay disparities based on gender/race). We consider a subset of 4 features (race, sex, education, age) and one response variable $Y$ (income). The favorable outcome for $Y$ is $Y\geq\$50$. The protected attributes are $G$ = {race, sex} with “white” and “male” being the privileged groups, respectively. These variables were preprocessed as follows. All non-white races were collapsed into one “non-white” category, age was compressed into decades, and education below 11th grade and above a Bachelor’s degree were combined into one category each. Table 1 summarizes the variables.
@@
Age (years) {17-26, 27-36, …, 87-96}
Education {< 11th grade, 11th grade, High school, Some college,
@@
[…]
Table 1: Variables in the experiment on the Adult dataset.
The COMPAS dataset contains criminal history information for 6,172 defendants. While it is more commonly used in fairness literature, it also carries privacy concerns; even inferring membership in the dataset is equivalent to disclosing that an individual was accused of a crime. We consider a subset of five feature variables (race, sex, age, number of priors, and degree of charge) and one response variable $Y$ (recidivism within two years) on a subset of 5,278 “African-American” and “Caucasian” individuals. The favorable outcome for $Y$ is that the individual did not recidivate. The protected attributes are $G$ = {race, sex} with “Caucasian” and “female” being the privileged groups, respectively. Table 2 summarizes the variables. Similarly, we discretized age and number of priors into 3 categories.
@@
Table 2: Variables in the experiment on the COMPAS dataset.
5.2 Experiment and Algorithmic settings
For the AIM method, we set the workload of marginals $W$ to be all two-way marginals in both experiments. Though Algorithm 2 uses $\rho$-zCDP, our results are presented using $(\varepsilon,\delta)$-DP with $\delta=10^{-9}$ (which is $o(1/n)$) via the conversion theorem. This facilitates comparisons with other works, as $(\varepsilon,\delta)$-DP is the most common type of DP used in the literature.
In the Adult experiment, we set $\phi_{A}(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=$ 3 in equations (9) and (10) if education changes by more than one stage or if age changes by more than a decade, 2 if age changes by a decade, 1 if income decreases, and 0 otherwise. If more than one condition is satisfied, $\phi_{A}$ outputs the largest possible result.
In the COMPAS experiment, we define individual distortion functions $\phi_{i}$ for each variable (e.g., $\phi_{\textnormal{Age}}$) that all output 0 for no change. Additionally, $\phi_{\textnormal{Age}}$ and $\phi_{\textnormal{Priors}}$ output 1 if the value is changed to the adjacent category and 2 if changed to a non-adjacent category (e.g., < 25 to > 45 for age). $\phi_{\textnormal{Charge}}$ outputs 1 if the value is changed from felony to misdemeanor and 2 if changed from misdemeanor to felony. $\phi_{\textnormal{Recidivism}}$ outputs 2 if the value is changed. We define the distortion function $\phi_{C}$ for equations (9) and (10) as the sum of the individual distortion functions.
In both experiments, we convert $\phi_{A}$ and $\phi_{C}$ into a binary $\phi$ using threshold values $(0.99,1.99,2.99)$. We set $\mathbf{c}=(0.1,0.05,0)$ corresponding to each threshold for the distortion condition (equation (10)), so that a few minor changes to each record are permissible with small probability, but large changes or changes to too many attributes are strongly discouraged. For example, in the Adult experiment, the learned fairness transformation $q$ decreases income with probability $0.1$ and changes age by at most a decade with probability $0.05$, but is not allowed to change education by more than one stage or age by more than one decade. Any other change is freely permitted. We also fix each $\eta_{y,g_{1},g_{2}}=\eta$ for a given value $\eta$.
In each experiment, we split the original data into a training set and a test set in an 80/20 ratio and encode the favorable result as 1 and the unfavorable result as 0. Synthetic data of the same sample size as the training data were generated via SAFES in Algorithm 1 using AIM as the DP data synthesizer and TOT for the fairness transformation. As the privacy- and fairness- aware transformations in SAFES are probabilistic, we run 35 repeats to summarize the average performance and stability. These repeats are performed at a list of different privacy budget parameter $\varepsilon$ and fairness parameter $\eta$ described in Table 3 (exception for the original dataset). The values of $\eta_{1}$ and $\eta_{2}$ were chosen to have some separation in the metric values among different settings. Other values could just as easily have been chosen, and we would expect a continuous transition of metric values between chosen $\boldsymbol{\eta}$ (i.e., $\eta=0.05$ would produce a line in Figure 3 somewhere in between the $\eta=0.025$ and $\eta=0.1$ lines).
@@ 5.2 Experiment and Algorithmic settings
fairness $\eta\in\{\eta_{1},\eta_{2}\}$ fairness-aware only SAFES
Table 3: Privacy and fairness algorithmic settings. $(\eta_{1},\eta_{2})=(0.025,0.1)$ for Adult, $(0.08,0.15)$ for COMPAS. The workload $W$ in the AIM synthesizer contained all two-way marginals in both experiments.
We use the SmartNoise SDK (OpenDP,, 2023) for the AIM subroutine in the experiments, and the AIF360 library (Bellamy et al.,, 2018) for the TOT subroutine. The randomness associated with DP occasionally makes the fairness transformation infeasible for small $\varepsilon$ (e.g., $\approx 30\%$ of the attempts failed for $\varepsilon\in\{10^{-2},10^{-1.5}\}$ and $\eta=0.08$ for COMPAS). Therefore, even though 35 repeats were attempted for each combination of $\varepsilon$ and $\eta$, a small number of the results presented in Figures 3 and 4 are summarized based on $<35$ repeats.
To evaluate the utility of the synthetic data, we 1) compare the data with the original training data by measuring the TVD between all one-, two-, and three-way marginals for the datasets; 2) perform a Kolmogorov-Smirnov (KS) test to evaluate the distributional similarity between the original training and the synthetic datasets; and 3) measure, for a logistic regression classifier trained on the synthetic data, the prediction accuracy, false positive (FP) and false negative (FN) rates, F1-score, and the ROC AUC on the test data.
To evaluate fairness, we measure the COD of the synthetic data (to understand the change in structural bias) and the SPD, the AOC, and the FP and FN rate balance for the logistic classifier fairness. Each of these metrics are evaluated with the protected attribute being race and sex individually, as well as jointly (e.g., white male as the privileged group).
5.3 Results
Examples of graphical models obtained via Algorithm 2 for the Adult and COMPAS experiments are shown in Figure 2. All the general utility, classification performance, and fairness metrics yield similar insights. For this reason, as well as to conserve space, we present results for a representative sample of utility and fairness metrics in Figures 3 and 4. Complete numeric results and additional figures are found in the supplementary information.

### v3

@@ 3.3 Fairness-aware data preprocessing
for all $y$ and groups $g_{1},g_{2}\in G$. $\eta$ controls the trade-off between enforcing group fairness and maintaining statistical relationships in the input dataset; different $\eta$ can be used for different groups or response values. We denote $\eta_{y,g_{1},g_{2}}$ for different $\{y,g_{1},g_{2}\}$ collectively by $\boldsymbol{\eta}$.
[…]
@@
$P\left(\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=1|G=\mathbf{g},X=\mathbf{x},Y=y\right)\leq c_{\mathbf{g},\mathbf{x},y},$ (10)
where 1 indicates the largest possible distortion – an individual jumps from one group in the original to the other one after the transformation. A small $c_{\mathbf{g},\mathbf{x},y}$ value would limit the probability of that happening. Similar to $\eta$, different values of $c_{\mathbf{g},\mathbf{x},y}\in[0,1]$ may be specified for different groups or response values. We denote the bounds $c_{\mathbf{g},\mathbf{x},y}$ collectively as $\mathbf{c}$. The distortion function $\phi$ is also customizable based on the context of the problem, and there is not necessarily a single best function for a given case. For example, $\phi$ that modifies an individual’s age by several decades should output a high distortion value $\phi$, while $\phi$ that does not change an individual’s age would output a 0 distortion value.
Finally, to ensure the transformation does not drastically modify the distribution of the input dataset, TOT minimizes the total variation distance (TVD) as the loss function, which, for discrete attributes, is
@@
Equations (8), (9), and (11), plus the constraint that $q_{\tilde{X},\tilde{Y}|D}$ needs to a proper distribution, give the optimization
@@
$J(q_{\tilde{Y}|G}(y|g_{1}),q_{\tilde{Y}|G}(y|g_{2}))\leq\eta_{y,g_{1},g_{2}},$
$\mathbb{E}\left[\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})|G=\mathbf{g},X=\mathbf{x},Y=y\right]\leq c_{\mathbf{g},\mathbf{x},y}.$
Equation (12) can be solved via a standard convex solver like the embedded conic solver (ECOS) (Domahidi et al.,, 2013) available in the Python CVXPY library (Diamond and Boyd,, 2016).
The RW method
RW Kamiran et al., (2012) is a fairness data preprocessing method that debias the data distribution for specific ML tasks. Unlike TOT, which learns a probabilistic transformation of $(X,Y)$, RW leaves $D=(X,Y,G)$ unchanged but assigns a weight to each record in $D$ so that, under the weighted empirical distribution, the sensitive attribute $G$ and the label $Y$ are approximately independent. Let $\hat{P}$ denote the empirical probability and define
$P(G=g,Y=c)\;\triangleq\;\hat{P}(G=g)\,\hat{P}(Y=c),$ (13)
implying $G$ and $Y$ are independent under equation (13). RW designs instance-level reweighting schemes so that $G$ and $Y$ are independent (i.e., equation (13) holds) after the procedures. One such scheme is
$W(g,c)\;=\;\frac{P(G=g,Y=c)}{\hat{P}(G=g,Y=c)}\,.$ (14)
Define the reweighted joint distribution $\hat{P}_{\,W}(G=g,Y=c)\;=\;W(g,c)\,\hat{P}(G=g,Y=c),$ which is equal to $P(G=g,Y=c)$ equation (13), ensuring $G$ and $Y$ is independent after the reweighting.
Classification models (e.g., weighted logistic regression or support vector machine) built upon the reweighted samples enjoy improved fairness. RW is parameter-free, fast to compute, and does not re-assign individuals to different groups or relabel individuals with different outcomes.
@@ 4 Experiments
We run experiments using the SAFES procedure on three real-world datasets: Adult (Becker and Kohavi,, 1996), the Correctional Offender Management Profiling for Alternative Sanctions (COMPAS) (ProPublica,, 2016), and German Credit Hofmann, (1994) data. Each dataset is publicly available for download. We evaluate the three-way privacy-fairness-utility trade-offs in SAFES against data synthesis or preprocessing methods that focus on either privacy preservation or fairness enhance, but not both. We also compare the results to those obtained when the original dataset, rather than a synthetic counterpart, is used directly (i.e., no privacy or fairness constraints, but still undergoing the preprocessed feature sets described in Section 4.1) so as to directly separate and understand the privacy-fairness-utility trade-off as a function of $\varepsilon$ and $\eta$.
To evaluate the general utility of the synthetic data via SAFES, we 1) compare the synthetic data with the original data by measuring the TVD between all one-, two-, and three-way marginals for the datasets and 2) perform a Kolmogorov-Smirnov (KS) test to evaluate the distributional similarity between the original and the synthetic datasets. TO evaluate ML utility of the synthetic data, We train a classifier on the synthetic data, and calculate the prediction accuracy, false positive (FP) and false negative (FN) rates, F1 score, and the ROC AUC on a test data. To evaluate fairness, we measure the COD of the synthetic data, the SPD, the AOC, and the FP and FN rate balance for the fairness of the trained classifier.
@@ 4.1 Data
We use the Adult data in three types of experiments: evaluating SAFES(AIM + TOT) for synthetic data utility and fairness when the original data $D$ contains only categorical variables (Section 4.2), evaluating SAFES(DP-CTGAN + RW)for synthetic data utility and fairness when $D$ contains a mixture of categorical and continuous variable (Section 4.3), and sensitivity analysis of the synthetic data to hyper-parameter specification (Section 4.4). The COMPAS dataset is used for the SAFES(AIM + TOT) experiment in Section 4.2. We use the German credit dataset to demonstrate SAFES(DP-CTGAN + RW) on a dataset with a larger number of mixed categorical and continuous variables.
Adult
The Adult dataset is a subset of US 1994 Census income data on 48,842 individuals. It is frequently used in both privacy and fairness literature due to the inclusion of sensitive variables (e.g., income) and the encoding of real-world discrimination (e.g., pay disparities based on gender/race). We consider a subset of 4 features (race, sex, education, age) and one response variable $Y$ (income). The favorable outcome for $Y$ is \$50,000. The protected attributes are $G$ = {race, sex} with white and male being the privileged groups, respectively. These variables were preprocessed as follows. All non-white races were collapsed into one non-white category; for the experiment in Sections 4.2 and 4.4, age was compressed into decades; for the experiment in Section 4.3, it was kept as is; education below 11th grade and above a Bachelor’s degree were combined into one category each. Table 1 summarizes the variables.
Table 1: Variables in the experiments on the Adult dataset.
@@
Age (years) Categorical: {17-26, 27-36, …, 87-96}; Continuous: [17, 96]
Education {<<11th grade, 11th grade, High school, Some college,
@@
Income {>>\$50k, $\leq$\$50k}
COMPAS
This dataset contains criminal history information for 6,172 defendants. While it is more commonly used in fairness literature, it also carries privacy concerns – even inferring membership in the dataset is equivalent to disclosing that an individual was accused of a crime. We consider a subset of five feature variables (race, sex, age, number of priors, and degree of charge) and one response variable $Y$ (recidivism within two years) on a subset of 5,278 African-American and Caucasian individuals. The favorable outcome for $Y$ is that the individual did not recidivate. The protected attributes are $G$ = {race, sex} with Caucasian and female being the privileged groups, respectively. We discretized age and number of priors into 3 categories. Table 2 summarizes the variables.
Table 2: Variables in the experiment on the COMPAS dataset.
@@
German Credit
This dataset contains 1000 loan applications and is widely used in fairness literature. It has 20 categorical and numerical attributes. The response variable $Y$ is binary, indicating creditworthiness (Good vs. Bad). We take the protected attribute as sex, derived from the personal status and sex field by mapping {male: divorced/separated, married/widowed, single} to Male and {female: divorced/separated/married, single} to Female and treat Male as the privileged group. As a preprocessing step, we rescale the attribute credit amount by dividing it with $1000$. Table 9 in the Appendix summarizes the variables.
4.2 SAFES (AIM + TOT) for categorical data
We use the Adult (discretized age) and the COMPAS data as listed in Table 1 and 2 for this experiment.
Privacy settings
We set the workload $W$ in the AIM Algorithm 2 to be all two-way marginals in both datasets. We pre-specify the overall privacy loss in term of $(\varepsilon,\delta)$-DP with a range of $\varepsilon$ given in Table 3 and $\delta=10^{-9}=o(1/n)$.
Table 3: Privacy and fairness algorithmic settings. SAFES is compared to the other three cells on utility and fairness metrics.
@@ 4.2 SAFES (AIM + TOT) for categorical data
fairness $\eta\in\{0.025,0.1\}$ for Adult; fairness-aware only SAFES
{0.08, 0.15} for COMPAS
Fairness settings
We examine two fairness parameter $\eta$ values in each of the two datasets as listed in Table 3. The values were selected to create separation in the fairness metrics, allowing us to visualize how the metrics changes as $\eta$ varies. Other $\eta$ values could have been used and we would expect a monotonic transition of the metric values beyond or between the results obtained at our examined $\boldsymbol{\eta}$ values. In the Adult data, we set $\phi(\mathbf{x},y,\tilde{\mathbf{x}},\tilde{y})=3$ if education changes by more than one stage or if age changes by more than a decade, 2 if age changes by a decade, 1 if income decreases, and 0 otherwise. In the COMPAS data, we set $\phi_{\textnormal{Age}}$ and $\phi_{\textnormal{Priors}}$ to be 1 if the value is changed to the adjacent category and 2 if changed to a non-adjacent category (e.g., ¡ 25 to ¿ 45 for age); $\phi_{\textnormal{Charge}}$ outputs 1 if the value is changed from felony to misdemeanor and 2 if changed from misdemeanor to felony; $\phi_{\textnormal{Recidivism}}$ outputs 2 if the value is changed; $\phi$ for each variable output 0 if there is no change. For coding and interpretability reasons, in both datasets, we convert the $\phi$ functions described above into a binary $\phi$ as defined in equation (10) with threshold values $(0.99,1.99,2.99)$ and $\mathbf{c}=(0.1,0.05,0)$ corresponding to each threshold. The output values from $\phi$ are selected in accordance with common sense and societal norms, though alternative values can be adopted when suitable. The small $\mathbf{c}$ values implies that only minor changes to each record are permissible with small probability and large changes or changes to too many attributes are strongly discouraged. For example, in the Adult data, the fairness preprocessing decreases income with probability $0.1$ and changes age by at most a decade with probability $0.05$, but is not allowed to change education by more than one stage or age by more than one decade. Any other change is freely permitted. We evaluate the sensitivity of the fairness-utility trade-off to different choices of $\mathbf{c}$ in Section 4.4.
Experiment settings
For both the Adult and COMPAS data, we split the original dataset (after the preprocessing described in Section 4.1) into a training set and a test set in an 75/25 ratio. Using the training data as input, synthetic data of the same sample size as the training data were generated via SAFES in Algorithm 1 using AIM as the DP data synthesizer and TOT for the fairness transformation. As the privacy- and fairness- aware transformations in SAFES are probabilistic, we run 35 repeats to summarize the average performance and stability, for each of combination settings defined by the privacy loss parameters $\varepsilon$ and fairness parameters $\eta$ listed in Table 3. We use the SmartNoise SDK (OpenDP,, 2023) for the AIM subroutine in the experiments, and the AIF360 library (Bellamy et al.,, 2018) for the TOT subroutine. The randomness associated with DP occasionally makes the fairness transformation infeasible for small $\varepsilon$ (e.g., $\approx 30\%$ of the attempts failed for $\varepsilon\in\{10^{-2},10^{-1.5}\}$ and $\eta=0.08$ for COMPAS). Therefore, even though 35 repeats were attempted for each combination of $\varepsilon$ and $\eta$, a small number of the results presented below are summarized based on <35<35 repeats. For the downstream classifier trained on the SAFES synthetic data, we use logistic regression. The fairness metrics are evaluated with the protected attribute being race and sex individually, as well as jointly (e.g., white male as the privileged group).
Results
Examination of the general utility, classification performance, and fairness metrics suggests similar insights across the board. For this reason and to conserve space, we present the results for a representative subset of the utility and fairness metrics in Figs. 2 and 3. Complete tabular numeric results and additional figures are found in the Appendix. The observations from the figures on the privacy-fairness-utility trade-off are summarized below. The findings in (1) to (3) below can also be observed in Figs. 1 - 14 in the Appendix.
  1. (1)
    As expected, stringent privacy (small $\varepsilon$) and fairness (small $\eta$) parameters result in higher general utility loss (larger marginal TVD and KS statistic) and worse classifier prediction performance (accuracy, F1 score, AUC, etc.). Owing to its dual focus on privacy preservation and fairness improvement, SAFES-generated synthetic data tends to be inferior to privacy-only and fairness-only synthetic data in the utility metrics, following the expected three-way privacy-fairness-utility trade-off. However, both experiments suggest that similar-to-baseline utility can be achieved with improved fairness metrics for $\varepsilon\leq 1$ (generally considered strong privacy), demonstrating the value of SAFES compared to privacy- or fairness-only approaches.
  2. (2)
    The performance of the classifiers trained on SAFES synthetic data is similar to that obtained in other works on DP data synthesis using these datasets. For example, the $\approx 0.8$ accuracy in the Adult experiment (Figure 3 in the Appendix) matches the accuracy obtained by Pujol et al., (2023); the $\approx 0.7$ AUC in the COMPAS experiment (Figure 10 in the Appendix) outperforms the AUC by Pereira et al., (2024).
  3. (3)
    DP guarantees at small $\varepsilon$ is more influential on general utility than the fairness enhancement quantified by $\eta$. For larger $\varepsilon$, $\eta$ has a more noticeable impact on general utility. The fairness transformation generally has little impact on the classifier prediction performance, with or without DP guarantees.
  4. (4)
    Even for strong privacy guarantees and fairness constraints ($\varepsilon\in(0.05,1)$ and small $\eta$), synthetic data via SAFES have similar utility to the original by several metrics, including the TVD of three-way marginals.
  5. (5)
    In general, at very small $\varepsilon$, the SAFES synthetic data and the DP-only synthetic data display more fairness than the original. Another way to interpret this is that the SAFES synthetic data do not outperform the DP-only synthetic data in the fairness metrics at small privacy budgets. This is due to the large amount of noise injected for the DP guarantees, but comes at a cost of poor utility. For $\varepsilon$ large enough to avoid this but still small enough to yield strong privacy guarantees, the synthetic data show significantly enhanced fairness for all three protected groups compared to the original that favors the privileged group, along with similar utility to the original.
  6. (6)
    Fairness guaranteed by SAFES at small $\eta$ is robust to changes in privacy loss and stabilizes as $\eta$ decreases, allowing confident calibration of the privacy-utility balance while maintaining fairness.
Figure 2: Examples of the privacy (points on each line) vs fairness (y-axis) vs utility (x-axis) trade-off in the Adult experiment. In each plot, each point on a line represents the mean and the error bar indicates $\pm 1$ SD over 35 repeats at a different privacy loss $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; lines represent different fairness parameters $\eta$; $x$-axis values further left correspond to better utility.
Figure 3: Examples of the privacy (points on each line) vs fairness (y-axis) vs utility (x-axis) trade-off in the COMPAS experiment. In each plot, each point on a line represents the mean and the error bar indicates $\pm 1$ SD over 35 repeats at a different privacy loss parameter $\varepsilon$ value $\in\{10^{-2}\;(\mbox{rightmost}),10^{-1.5},10^{-1},\ldots,10\;(\mbox{leftmost})\}$; lines represent different fairness parameters $\eta$; $x$-axis values further left correspond to better utility.
Figs. 2 and 3 suggest the FN rate balance in several cases is positive rather than negative. At first glance, this seems to indicate that the unprivileged sex group (female in Adult, male in COMPAS) is more likely to have the favorable outcome per this metric. However, a positive value is rather an indication that more unprivileged than privileged individuals are incorrectly classified to the unfavorable outcome, which is still a disadvantageous result for the unprivileged group. Additionally, when measuring COD jointly on race+sex in the COMPAS experiment, the COD of the DP-only synthetic data (solid blue line) does not appear to converge to the COD of the original dataset (solid blue dot). This is an artifact of the graphical models constructed by the AIM subroutine in SAFES when selecting from only one- and two-way marginals, as we did in our evaluations. Three-way relationships, if they exist, such as between race, sex, and recidivism, may not be fully captured. Fig. 4 shows an example of this for the Adult and COMPAS experiments.